# Interpolation theorems for conjugations and applications

### v1

@@ 3.1. Main results
(iii)$\implies$(i). Let $V$ be an anti-unitary operator on $H$ such that $VNV^{-1}=-N^{*}$. By normality of $N$, there exists a conjugation $J$ on $H$ that satisfies $JNJ=N^{*}$. Hence, we have $VNV^{-1}=-JNJ$, and so, since $J^{2}=I$, we get
$(JV)N(V^{-1}J)=-J^{2}NJ^{2}=-N.$
Put $U=JV$; then $U\in{\mathcal{B}}(H)$ is unitary and $U^{*}=U^{-1}=(JV)^{-1}=V^{-1}J^{-1}=V^{-1}J$. Therefore, $UNU^{*}=-N$.
@@
(i)$\implies$(ii). Let $U\in{\mathcal{B}}(H)$ be a unitary operator satisfying $UN=-NU$. Fix a Borel subset $\Delta_{0}\subset\mathbb{C}$ such that $\mathbb{C}\setminus\Delta_{0}=-\Delta_{0}\cup\{0\}$. By Lemma 2.11, we have $UE_{N}(\Delta_{0})=E_{N}(-\Delta_{0})U$ and $UE_{N}(-\Delta_{0})=E_{N}(\Delta_{0})U$. Consequently,
$U{\rm ran}(E_{N}(\Delta_{0}))={\rm ran}(E_{N}(-\Delta_{0}))\quad\mbox{and}\quad U{\rm ran}(E_{N}(-\Delta_{0}))={\rm ran}(E_{N}(\Delta_{0})).$
@@
Given that $\mathbb{C}=\Delta_{0}\cup(-\Delta)\cup\{0\}$ is a partition, it follows from the above equalities that $U$ has the form
$U=\begin{bmatrix}0&*&0\\ W&0&0\\ 0&0&*\end{bmatrix}\begin{array}[]{l}{\rm ran}(E_{N}(\Delta_{0}))\\ {\rm ran}(E_{N}(-\Delta_{0}))\\ {\rm ran}(E_{N}(\{0\}))\end{array}$
Where $W$ is an isometric isomorphism between Hilbert spaces. Noting that $N$ can also be written as
$N=\begin{bmatrix}N_{1}&0&0\\ 0&N_{2}&0\\ 0&0&0\end{bmatrix}\begin{array}[]{l}{\rm ran}(E_{N}(\Delta_{0}))\\ {\rm ran}(E_{N}(-\Delta_{0}))\\ {\rm ran}(E_{N}(\{0\}))\end{array},$
and since $UN=-NU$, simple matrix calculations show that $WN_{1}=-N_{2}W$. This implies that $N_{1}\cong-N_{2}$, and hence $N_{|{\rm ker}(N)^{\perp}}=N_{1}\oplus N_{2}\cong N_{1}\oplus-N_{1}$. Therefore, $N$ is skew symmetric by [17, Theorem 1.10]. ∎
@@
Note that every bounded linear $T$ acting on a Hilbert space can be written as $T=N\oplus A$ where $N$ is normal and $A$ is an operator that has no reducing subspace for which the restriction of $A$ is normal (of course, either $N$ or $A$ could be absent). In fact, $N$ is the restriction of $T$ to the reducing subspace
(3.5) $H_{\rm nor}=\bigcap_{1\leq n,m<\infty}{\rm ker}({T^{*}}^{n}T^{m}-T^{m}{T^{*}}^{n}).$
@@
The following lemma is an immediate consequence of [28, Proposition 3.10] and (3.6).
@@
Let $M\in{\mathcal{B}}(H)$ be normal and $T\in{\mathcal{B}}(K)$ such that $M\oplus T$ is skew-symmetric. Then, $M\oplus T_{\rm nor}$ and $T_{\rm abnor}$ are skew-symmetric.
@@
Let $N\in{\mathcal{B}}(H)$ be a skew-symmetric normal operator and $\Delta\subseteq\mathbb{C}$ a Borel subset such $\Delta=-\Delta$. Then, the restriction of $N$ to the reducing subspace ${\rm ran}(E_{N}(\Delta))$ is skew-symmetric.
@@
Let $C$ be a conjugation on $H$ such that $CNC=-N^{*}$. By Lemma 2.11, we have $CE_{N}(\Delta)=E_{N}(-\Delta)C=E_{N}(\Delta)C$; hence, $C{\rm ran}(E_{N}(\Delta))={\rm ran}(E_{N}(\Delta))$. Let $C_{0}$ and $N_{0}$ denote the restrictions of $C$ and $N$ to ${\rm ran}(E_{N}(\Delta))$, respectively. Then, $C_{0}$ is a conjugation on $E_{N}(\Delta)$. Moreover, it follows immediately from the equality $CNC=-N^{*}$ that $C_{0}N_{0}C_{0}=-N_{0}^{*}$. ∎
In [16, Lemma 3.2] and [2, Corollary 3.4], the authors proved that if $N$ is normal, then a direct sum $N\oplus T$ is complex symmetric (resp. conjugate-normal) if and only if $T$ is complex-symmetric (resp. conjugate-normal). However, this property does not hold for skew-symmetry even if $N$ is skew-symmetric. The next result provides a complete description of normal operators $N$ that satisfy this property.
@@
Let $N\in{\mathcal{B}}(H)$ be normal. The followings statements are equivalents:
@@
    For every separable complex Hilbert space $K$ and every $T\in\mathcal{B}(K)$,
@@
(i)$\implies$(ii). Suppose statement (i) holds. Taking $T=0$, the zero operator on any Hilbert space $K$, we obtain that $N\oplus 0$ is skew-symmetric. By [17, Lemma 2.1], it follows that $N_{|({\rm ker}N)^{\perp}}$ is skew-symmetric. Therefore, $N=N_{|({\rm ker}N)^{\perp}}\oplus 0$ is skew-symmetric.
@@
In view of Lemma 3.6 and [17, Lemma 2.1], we can assume without loss of generality that $T$ is normal and $N=N_{|{\rm ker}(N)^{\perp}}$. Moreover, by (3.1) and (3.3), we may directly assume that
$N=\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}\quad\mbox{and}\quad H=\bigoplus_{1\leq i<\infty}L^{2}(\mu_{i})^{(i)}$
@@
Let $\Delta_{0}=\mathbb{C}\setminus\cup_{i\in\mathbb{Z}^{+}}\Delta_{i}$; clearly, $\Delta_{0}=-\Delta_{0}$. For convenience, put $N_{i}=M_{\mu_{i}}^{(i)}$, $H_{i}=L^{2}(\mu_{i})^{(i)}$ for $i\in\mathbb{Z}^{+}$, and let $T=T_{0}\oplus\bigoplus_{i\in\mathbb{Z}^{+}}T_{i}$ be the decomposition of $T$ with respect to the orthogonal decomposition $K={\rm ran}(E_{T}(\Delta_{0}))\oplus\bigoplus_{i\in\mathbb{Z}^{+}}{\rm ran}(E_{T}(\Delta_{i}))$. For $0\leq i<\infty$, put $K_{i}={\rm ran}(E_{T}(\Delta_{i}))$.
@@
${\rm ran}(E_{N\oplus T}(\Delta_{0}))=K_{0}\quad\mbox{and}\quad{\rm ran}(E_{N\oplus T}(\Delta_{i}))=H_{i}\oplus K_{i}\;\;\mbox{for every }i\in\mathbb{Z}^{+}.$
@@
Fix $i\in\mathbb{Z}^{+}$, and let us prove that $T_{i}$ is skew-symmetric. Given $\mu_{i}(\Delta)=\mu_{i}(-\Delta)$ for every Borel subset $\Delta\subseteq\mathbb{C}$, [17, Theorem 1.11] implies that $N_{i}$ is skew-symmetric; hence, $N_{i}\cong-N_{i}$ by Lemma 3.5. Combining this with (3.8), we obtain $N_{i}\oplus T_{i}\cong N_{i}\oplus-T_{i}$. Since $M_{\mu_{i}}$ is a star-cyclic normal operator ([8, Page 269, Theorem 3.4]), applying [8, Page 295, Proposition 10.6] $i$ times to $M_{\mu_{i}}$, we conclude that $T_{i}\cong-T_{i}$, and so $T_{i}$ is skew-symmetric by Lemma 3.5.
@@
Let $N\in{\mathcal{B}}(H)$ be normal, and suppose that there exist reducing subspaces $\{H_{i}\}_{1\leq i\leq r}$, $r<\infty$, of $N$ such that $H=\vee(\sum_{i=1}^{r}H_{i})$ and $N_{|H_{i}}$ is star-cyclic for each $i\in\{1,2,\dots,r\}$. Then, for every reducing subspace $L$ of $N$, we have $m(N_{|L})<\infty$.
@@
First, let us show that $N$ is a finite direct sum of star-cyclic normal operators. We will proceed by induction on $r$. The result is clear when $r=1$. Assume the lemma holds for some integer $r\geq 1$. Let $H_{1},H_{2},\dots H_{r+1}$ be $r+1$ reducing subspaces of $N$ such that $H=\vee(\sum_{i=1}^{r+1}H_{i})$ and $N_{|H_{i}}$ is star cyclic for $i\in\{1,2,\dots,r+1\}$. Let $K=\vee(\sum_{i=1}^{r}H_{i})$ and $P\in{\mathcal{B}}(H)$ be the orthogonal projection onto $K$. Then, $K$ reduces $N$; moreover, by the induction hypothesis, $N_{|K}$ is a finite direct sum of star-cyclic operators.
By star-cyclicity of $N_{|H_{r+1}}$, there exists $h\in H$ such that $H_{r+1}=L_{N,h}$. To complete the proof, it is clear that it suffices to show that $H=K\oplus L_{N,(I-P)h}$. Note that $K$ and $L_{N,(I-P)h}$ are orthogonal. In fact, since $K$ reduces $N$, its orthogonal projection $P$, and hence $(I-P)$, commutes with $N$ and $N^{*}$. Consequently,
$L_{N,(I-P)h}$ =$=\bigvee\big{\{}p(N,N^{*})(I-P)h:p\in\mathbb{C}[x,y]\big{\}}$
$=(I-P)\bigvee\big{\{}p(N,N^{*})h:p\in\mathbb{C}[x,y]\big{\}}\subseteq K^{\perp}.$
It remains to prove that $H=K+L_{N,(I-P)h}$; we need only to show that $H_{r+1}\subseteq K+L_{N,(I-P)h}$. For $p\in\mathbb{C}[x,y]$, we have $p(N,N^{*})h=p(N,N^{*})Ph+p(N,N^{*})(I-P)h$. Since $Ph\in K$ and $K$ reduces $p(N,N^{*})$, we have $p(N,N^{*})Ph\in K$. Hence, $p(N,N^{*})h\in K+L_{N,(I-P)h}$, and thus $H_{r+1}\subseteq K+L_{N,(I-P)h}$.
@@
Since $N$ is a direct sum of $n$ star-cyclic normal operators $M_{1},M_{2},\dots,M_{n}$ with $n<\infty$, applying [8, Page 295, Proposition 10.6] to each $M_{i}$, we conclude $M_{\mu}\cong M_{\mu}^{(2)}$; which contradicts (3.2). ∎
@@
(ii)$\implies$(iii). First, let us show that $N_{|L_{N,Z}^{\perp}}$ is skew-symmetric. By the argument in (2.7), there exists a conjugation $J$ on $H$ so that $(JV)N=-N(JV)$ and $JN^{*}J=N$. By the Fuglede-Putnam Theorem, it follows that $(JV)N^{*}=-N^{*}(JV)$; hence, $VN^{*}=J^{2}VN^{*}=-JN^{*}JV=-NV$. Consequently, for any polynomial $p\in\mathbb{C}[x,y]$, one can easily see that $Vp(N,N^{*})=p(-N,-N^{*})^{*}V$. Therefore, for every $i\in I$,
$Vp(N,N^{*})x_{i}=p(-N,-N^{*})^{*}Vx_{i}=p(-N,-N^{*})^{*}y_{i}$

### v3

@@ 3.1. Main results
(ii)$\implies$(iii). Let $V$ be an anti-unitary operator on $\mathcal{H}$ such that $VN_{k}V^{-1}=-N_{k}^{*}$ for every $k\in K$. As in (2.10), there exists a conjugation $J$ on $\mathcal{H}$ that satisfies $JN_{k}J=N_{k}^{*}$ for every $k\in K$. Hence, we have $VN_{k}V^{-1}=-JN_{k}J$, and so, since $J^{-1}=J$, we get
(3.7) $(JV)N_{k}(V^{-1}J)=-J^{2}N_{k}J^{2}=-N_{k}\quad\mbox{for every $k\in K$}.$
Put $U=JV$; then $U\in{\mathcal{B}}(\mathcal{H})$ is unitary and $U^{*}=U^{-1}=(JV)^{-1}=V^{-1}J^{-1}=V^{-1}J$. Therefore, for any $k\in K$, $UN_{k}U^{*}=-N_{k}$.
(iii)$\implies$(i). It is easy to see that we may assume without loss of generality that $K$ is infinite. Let $U\in{\mathcal{B}}(\mathcal{H})$ be a unitary operator satisfying $UN_{k}=-N_{k}U$ for every $k\in K$. First, let us reduce the problem to the case where $\bigcap_{k\in K}{\rm ker}(N_{k})=\{0\}$. Denote this subspace by $\mathcal{L}_{\infty}$. One can easily see that $\mathcal{L}_{\infty}$ reduces $U$, and that each $N_{k}$ has the form $N_{k}=M_{k}\oplus 0$ with respect to the decomposition $\mathcal{H}=\mathcal{L}_{\infty}^{\perp}\oplus\mathcal{L}_{\infty}$. Clearly, $\bigcap_{k\in K}{\rm ker}(M_{k})=\{0\}$; hence, if a conjugation $C$ on $\mathcal{L}_{\infty}^{\perp}$ exists such that $CM_{k}C=-M_{k}^{*}$ for each $k\in K$, then one can see that the map $C\oplus C_{\infty}$, where $C_{\infty}$ is an arbitrary conjugation on $\mathcal{L}_{\infty}$, is a conjugation on $\mathcal{H}$ that satisfies all the desired equalities.
Note that we can further assume that there exists a countable subset $\{N_{k_{p}}\}_{p\in\mathbb{Z}^{+}}$ of $\{N_{k}\}_{k\in K}$ such that $\bigcap_{p\in\mathbb{Z}^{+}}{\rm ker}(N_{k_{p}})=\{0\}$. In fact, since $\mathcal{H}=\bigvee\left(\bigcup_{k\in K}{\rm ker}(N_{k})^{\perp}\right)$, one can use the separability of $\mathcal{H}$ to show that there exists a countable subset $\{k_{p}\}_{p\in\mathbb{Z}^{+}}\subseteq K$ so that $\mathcal{H}=\bigvee\left(\bigcup_{p\in\mathbb{Z}^{+}}{\rm ker}(N_{k_{p}})^{\perp}\right)$.
@@
Fix a Borel subset $\Delta_{0}\subset\mathbb{C}$ such that $\mathbb{C}\setminus\Delta_{0}=-\Delta_{0}\cup\{0\}$. By Lemma 2.15, we have $UE_{N_{k_{1}}}(\Delta_{0})=E_{N_{k_{1}}}(-\Delta_{0})U$ and $UE_{N_{k_{1}}}(-\Delta_{0})=E_{N_{k_{1}}}(\Delta_{0})U.$ Consequently,
(3.8) $U{\rm ran}(E_{N_{k_{1}}}(\Delta_{0}))={\rm ran}(E_{N_{k_{1}}}(-\Delta_{0}))\;\;\mbox{and}\;\;U{\rm ran}(E_{N_{k_{1}}}(-\Delta_{0}))={\rm ran}(E_{N_{k_{1}}}(\Delta_{0})).$
@@
For convenience, put $\mathcal{H}_{+}={\rm ran}(E_{N_{k_{1}}}(\Delta_{0}))$, $\mathcal{H}_{-}={\rm ran}(E_{N_{k_{1}}}(-\Delta_{0}))$ and $\mathcal{L}_{1}={\rm ran}(E_{N_{k_{1}}}(\{0\}))={\rm ker}(N_{k_{1}})$. Given that $\mathbb{C}=\Delta_{0}\cup(-\Delta_{0})\cup\{0\}$ is a partition, it follows from (3.8) that $U$ has the form
$U=\begin{bmatrix}0&*&0\\ W&0&0\\ 0&0&U_{1}\end{bmatrix}\begin{array}[]{l}\mathcal{H}_{+}\\ \mathcal{H}_{-}\\ \mathcal{L}_{1}\end{array}$
where $W$ and $U_{1}$ are isometric isomorphisms between Hilbert spaces. Additionally, since for every $k\in K$,
$N_{k}E_{N_{k_{1}}}(\Delta_{0})=E_{N_{k_{1}}}(\Delta_{0})N_{k}\quad\mbox{and}\quad N_{k}E_{N_{k_{1}}}(-\Delta_{0})=E_{N_{k_{1}}}(-\Delta_{0})N_{k},$
we can also write
$N_{k}=\begin{bmatrix}N_{k,+}&0&0\\ 0&N_{k,-}&0\\ 0&0&R_{k,1}\end{bmatrix}\begin{array}[]{l}\mathcal{H}_{+}\\ \mathcal{H}_{-}\\ \mathcal{L}_{1}\end{array}$
where $N_{k,+}$, $N_{k,-}$ and $R_{k,1}$, $k\in K$, are normal operators such that the $R_{k,1}$’s are mutually commuting.
Since $UN_{k}=-N_{k}U$, simple matrix calculations show that
(3.9) $WN_{k,+}=-N_{k,-}W\quad\mbox{and}\quad U_{1}R_{k,1}=-R_{k,1}U_{1}\quad\mbox{for every $k\in K$}.$
Put $\mathcal{M}_{1}=\mathcal{H}_{+}\oplus\mathcal{H}_{-}$ and let $V:\mathcal{M}_{1}\to\mathcal{H}_{-}\oplus\mathcal{H}_{-}$ be the unitary operator given by
$\begin{array}[]{rcl}V:\mathcal{H}_{+}\oplus\mathcal{H}_{-}&\to&\mathcal{H}_{-}\oplus\mathcal{H}_{-}\\ (x,y)&\mapsto&(Wx,y).\end{array}$
Then, it is easy to see that $V{N_{k}}_{|\mathcal{M}_{1}}V^{-1}=(-N_{k,-})\oplus N_{k,-}$ for every $k\in K$. Since the normal operators $\{N_{k,-}\}_{k\in K}$ are mutually commuting, as in (2.10), there exists a conjugation $J$ on $\mathcal{H}_{-}$ such that $JN_{k,-}J=\left(N_{k,-}\right)^{*}$ for every $k\in K$. Put $J_{1}=\begin{bmatrix}0&J\\ J&0\\ \end{bmatrix}\begin{array}[]{l}\mathcal{H}_{-}\\ \mathcal{H}_{-}\end{array}.$ It is easy to see that $J_{1}$ is a conjugation on $\mathcal{H}_{-}\oplus\mathcal{H}_{-}$ satisfying
$J_{1}\left((-N_{k,-})\oplus N_{k,-}\right)J_{1}=-\left((-N_{k,-})\oplus N_{k,-}\right)^{*}\quad\mbox{for every $k\in K$}.$
Hence, $C_{1}:=V^{-1}J_{1}V$ is a conjugation on $\mathcal{M}_{1}$ that satisfies $C_{1}{N_{k}}_{|\mathcal{M}_{1}}C_{1}=-\left({N_{k}}_{|\mathcal{M}_{1}}\right)^{*}$ for every $k\in K$.
Since $\{R_{k,1}\}_{k\in K}$ are mutually commuting and $U_{1}R_{k,1}U_{1}^{*}=-R_{k,1}$ for every $k\in K$, by applying the preceding argument to $R_{k_{2},1}\in\mathcal{B}(\mathcal{L}_{1})$, we obtain subspaces $\mathcal{M}_{2}$ and $\mathcal{L}_{2}$ of $\mathcal{L}_{1}$, with $\mathcal{L}_{1}=\mathcal{M}_{2}\oplus\mathcal{L}_{2}$, a conjugation $C_{2}$ on $\mathcal{M}_{2}$ such that
$C_{2}{N_{k}}_{|\mathcal{M}_{2}}C_{2}=C_{2}{R_{k,1}}_{|\mathcal{M}_{2}}C_{2}=-\left({R_{k,1}}_{|\mathcal{M}_{2}}\right)^{*}=-\left({N_{k}}_{|\mathcal{M}_{2}}\right)^{*}\quad\mbox{for every $k\in K$},$
a unitary operator $U_{2}$ on $\mathcal{L}_{2}$, and a set of mutually commuting normal operators $\{R_{k,2}\}_{k\in K}$ on $\mathcal{L}_{2}$ such that ${N_{k}}_{|\mathcal{L}_{2}}=R_{k,2}$ and $U_{2}R_{k,2}U_{2}^{*}=-\left(R_{k,2}\right)^{*}$. This process produces a sequence $\mathcal{M}_{i}$, $i\geq 1$, of reducing subspaces of $N_{k}$, for $k\in K$, and conjugations $C_{i}$ on $\mathcal{M}_{i}$ such that ${N_{k}}_{|\mathcal{M}_{i}}$ is $C_{i}$-skew symmetric for all $i\in\mathbb{Z}^{+}$ and $k\in K$.
For each $i\geq 1$, it is easy to see that $\left(\sum_{j=1}^{i}\mathcal{M}_{j}\right)^{\perp}=\mathcal{L}_{i}=\bigcap_{j=1}^{i}{\rm ker}(N_{k_{j}}).$ Hence,
$\left(\bigoplus_{i=1}^{\infty}\mathcal{M}_{i}\right)^{\perp}=\bigcap_{i=1}^{\infty}\left(\sum_{j=1}^{i}\mathcal{M}_{j}\right)^{\perp}=\bigcap_{i=1}^{\infty}{\rm ker}(N_{k_{i}})=\{0\}.$
Consequently, $\mathcal{H}=\bigoplus_{i=1}^{\infty}\mathcal{M}_{i}$; moreover, one can check that $C:=\bigoplus_{i=1}^{\infty}C_{i}$ is a conjugation on $\mathcal{H}$ satisfying $CN_{k}C=-N_{k}^{*}$ for every $k\in K$. ∎
@@
Note that every bounded linear operator $T$ acting on a Hilbert space can be written as $T=N\oplus A$ where $N$ is normal and $A$ is an operator that has no reducing subspace for which the restriction of $A$ is normal (of course, either $N$ or $A$ could be absent). In fact, $N$ is the restriction of $T$ to the reducing subspace
(3.10) $\mathcal{H}_{\rm nor}=\bigcap_{1\leq n,m<\infty}{\rm ker}({T^{*}}^{n}T^{m}-T^{m}{T^{*}}^{n}).$
@@
The following lemma is an immediate consequence of [29, Proposition 3.10] and (3.11).
@@
Let $M\in{\mathcal{B}}(\mathcal{H})$ be normal and $T\in{\mathcal{B}}(\mathcal{K})$ such that $M\oplus T$ is skew-symmetric. Then, $M\oplus T_{\rm nor}$ and $T_{\rm abnor}$ are skew-symmetric.
@@
Let $N\in{\mathcal{B}}(\mathcal{H})$ be a skew-symmetric normal operator and $\Delta\subseteq\mathbb{C}$ a Borel subset such $\Delta=-\Delta$. Then, the restriction of $N$ to the reducing subspace ${\rm ran}(E_{N}(\Delta))$ is skew-symmetric.
@@
Let $C$ be a conjugation on $\mathcal{H}$ such that $CNC=-N^{*}$. By Lemma 2.15, we have $CE_{N}(\Delta)=E_{N}(-\Delta)C=E_{N}(\Delta)C$; hence, $C{\rm ran}(E_{N}(\Delta))={\rm ran}(E_{N}(\Delta))$. Let $C_{0}$ and $N_{0}$ denote the restrictions of $C$ and $N$ to ${\rm ran}(E_{N}(\Delta))$, respectively. Then, $C_{0}$ is a conjugation on $E_{N}(\Delta)$. Moreover, it follows immediately from the equality $CNC=-N^{*}$ that $C_{0}N_{0}C_{0}=-N_{0}^{*}$. ∎
In [17, Lemma 3.2] and [2, Corollary 3.4], the authors proved that if $N$ is normal, then a direct sum $N\oplus T$ is complex symmetric (resp. conjugate-normal) if and only if $T$ is complex-symmetric (resp. conjugate-normal). However, this property does not hold for skew-symmetry even if $N$ is skew-symmetric. The next result provides a complete description of normal operators $N$ that satisfy this property.
@@
Let $N\in{\mathcal{B}}(\mathcal{H})$ be normal. The followings statements are equivalent:
@@
    For every separable complex Hilbert space $\mathcal{K}$ and every $T\in\mathcal{B}(\mathcal{K})$,
@@
(i)$\implies$(ii). Suppose statement (i) holds. Taking $T=0$, the zero operator on any Hilbert space $\mathcal{K}$, we obtain that $N\oplus 0$ is skew-symmetric. By [18, Lemma 2.1], it follows that $N_{|({\rm ker}N)^{\perp}}$ is skew-symmetric. Therefore, $N=N_{|({\rm ker}N)^{\perp}}\oplus 0$ is skew-symmetric.
@@
In view of Lemma 3.8 and [18, Lemma 2.1], we can assume without loss of generality that $T$ is normal and $N=N_{|{\rm ker}(N)^{\perp}}$. Moreover, by (3.1) and (3.3), we may directly assume that
$N=\bigoplus_{1\leq i<\infty}M_{\mu_{i}}^{(i)}\quad\mbox{and}\quad\mathcal{H}=\bigoplus_{1\leq i<\infty}L^{2}(\mu_{i})^{(i)}$
@@
Let $\Delta_{0}=\mathbb{C}\setminus\cup_{i\in\mathbb{Z}^{+}}\Delta_{i}$; clearly, $\Delta_{0}=-\Delta_{0}$. For convenience, put $N_{i}=M_{\mu_{i}}^{(i)}$, $\mathcal{H}_{i}=L^{2}(\mu_{i})^{(i)}$ for $i\in\mathbb{Z}^{+}$, and let $T=T_{0}\oplus\bigoplus_{i\in\mathbb{Z}^{+}}T_{i}$ be the decomposition of $T$ with respect to the orthogonal decomposition $\mathcal{K}={\rm ran}(E_{T}(\Delta_{0}))\oplus\bigoplus_{i\in\mathbb{Z}^{+}}{\rm ran}(E_{T}(\Delta_{i}))$. For $0\leq i<\infty$, put $\mathcal{K}_{i}={\rm ran}(E_{T}(\Delta_{i}))$.
@@
${\rm ran}(E_{N\oplus T}(\Delta_{0}))=\mathcal{K}_{0}\quad\mbox{and}\quad{\rm ran}(E_{N\oplus T}(\Delta_{i}))=\mathcal{H}_{i}\oplus\mathcal{K}_{i}\;\;\mbox{forevery }i\in\mathbb{Z}^{+}.$
@@
Fix $i\in\mathbb{Z}^{+}$, and let us prove that $T_{i}$ is skew-symmetric. Given $\mu_{i}(\Delta)=\mu_{i}(-\Delta)$ for every Borel subset $\Delta\subseteq\mathbb{C}$, [18, Theorem 1.11] implies that $N_{i}$ is skew-symmetric; hence, $N_{i}\cong-N_{i}$ by Lemma 3.7. Combining this with (3.13), we obtain $N_{i}\oplus T_{i}\cong N_{i}\oplus-T_{i}$. Since $M_{\mu_{i}}$ is a star-cyclic normal operator ([9, p. 269, Theorem 3.4]), applying [9, p. 295, Proposition 10.6] $i$ times to $M_{\mu_{i}}$, we conclude that $T_{i}\cong-T_{i}$, and so $T_{i}$ is skew-symmetric by Lemma 3.7.
@@
Let $N\in{\mathcal{B}}(\mathcal{H})$ be normal, and suppose that there exist reducing subspaces $\{\mathcal{H}_{i}\}_{1\leq i\leq r}$, $r<\infty$, of $N$ such that $\mathcal{H}=\vee(\sum_{i=1}^{r}\mathcal{H}_{i})$ and $N_{|\mathcal{H}_{i}}$ is star-cyclic for each $i\in\{1,2,\dots,r\}$. Then, for every reducing subspace $L$ of $N$, we have $m(N_{|L})<\infty$.
@@
First, let us show that $N$ is a finite direct sum of star-cyclic normal operators. We will proceed by induction on $r$. The result is clear when $r=1$. Assume the lemma holds for some integer $r\geq 1$. Let $\mathcal{H}_{1},\mathcal{H}_{2},\dots\mathcal{H}_{r+1}$ be $r+1$ reducing subspaces of $N$ such that $\mathcal{H}=\vee(\sum_{i=1}^{r+1}\mathcal{H}_{i})$ and $N_{|\mathcal{H}_{i}}$ is star cyclic for $i\in\{1,2,\dots,r+1\}$. Let $\mathcal{K}=\vee(\sum_{i=1}^{r}\mathcal{H}_{i})$ and $P\in{\mathcal{B}}(\mathcal{H})$ be the orthogonal projection onto $\mathcal{K}$. Then, $\mathcal{K}$ reduces $N$; moreover, by the induction hypothesis, $N_{|\mathcal{K}}$ is a finite direct sum of star-cyclic operators.
By star-cyclicity of $N_{|\mathcal{H}_{r+1}}$, there exists $h\in\mathcal{H}$ such that $\mathcal{H}_{r+1}=L_{\{N\},\{h\}}$. To complete the proof, it is clear that it suffices to show that $\mathcal{H}=\mathcal{K}\oplus L_{\{N\},\{(I-P)h\}}$. Note that $\mathcal{K}$ and $L_{\{N\},\{(I-P)h\}}$ are orthogonal. In fact, since $\mathcal{K}$ reduces $N$, its orthogonal projection $P$, and hence $(I-P)$, commutes with $N$ and $N^{*}$. Consequently,
$L_{\{N\},\{(I-P)h\}}$ =$=\bigvee\big{\{}p(N,N^{*})(I-P)h:p\in\mathcal{P}_{2}(\mathbb{C})\big{\}}$
$=(I-P)\bigvee\big{\{}p(N,N^{*})h:p\in\mathcal{P}_{2}(\mathbb{C})\big{\}}\subseteq\mathcal{K}^{\perp}.$
It remains to prove that $\mathcal{H}=\mathcal{K}+L_{\{N\},\{(I-P)h\}}$; we need only to show that $\mathcal{H}_{r+1}\subseteq\mathcal{K}+L_{\{N\},\{(I-P)h\}}$. For $p\in\mathcal{P}_{2}(\mathbb{C})$, we have
$p(N,N^{*})h=p(N,N^{*})Ph+p(N,N^{*})(I-P)h.$
Since $Ph\in\mathcal{K}$ and $\mathcal{K}$ reduces $p(N,N^{*})$, we have $p(N,N^{*})Ph\in\mathcal{K}$. Hence, $p(N,N^{*})h\in\mathcal{K}+L_{\{N\},\{(I-P)h\}}$, and thus $\mathcal{H}_{r+1}\subseteq\mathcal{K}+L_{\{N\},\{(I-P)h\}}$.
@@
Since $N$ is a direct sum of $n$ star-cyclic normal operators $M_{1},M_{2},\dots,M_{n}$ with $n<\infty$, applying [9, p. 295, Proposition 10.6] to each $M_{i}$, we conclude $M_{\mu}\cong M_{\mu}^{(2)}$; which contradicts (3.2). ∎
@@
(ii)$\implies$(iii). First, let us show that $V_{1}$ exists. Since $VN_{k}V^{-1}=-N_{k}^{*}$, one can use (2.7) to show that $VN_{k}^{*}V^{-1}=-N_{k}$ for every $k\in K$. Consequently, for any $n\in\mathbb{Z}^{+}$, $p\in\mathcal{P}_{2n}(\mathbb{C})$ and $\{k_{l}\}_{l=1}^{n}\subseteq K$, one can easily see that
$Vp(N_{k_{1}},N_{k_{1}}^{*},\dots,N_{k_{n}},N_{k_{n}}^{*})=p(-N_{k_{1}},-N_{k_{1}}^{*},\dots,-N_{k_{n}},-N_{k_{n}}^{*})^{*}V.$
Therefore, for every $i\in I$,
$Vp(N_{k_{1}},N_{k_{1}}^{*},\dots,N_{k_{n}},N_{k_{n}}^{*})x_{i}$ $=p(-N_{k_{1}},-N_{k_{1}}^{*},\dots,-N_{k_{n}},-N_{k_{n}}^{*})^{*}Vx_{i}$
$=p(-N_{k_{1}},-N_{k_{1}}^{*},\dots,-N_{k_{n}},-N_{k_{n}}^{*})^{*}y_{i}$